If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
Answer by checkley71(8403)

LETS TAKE 2 AT A TIME
5*10/15=50/15=3.333
NOW COMBINE THIS WITH THE THE 15 MINUTE MIXER
3.33*15/3.33+15=49.95/18.33=2.73 MINUTES WHEN THEY WORK TOGETHER.
OR:
20/5=4 DRINKS PER MINUTE.
20/10=2 DRINKS PER MINUTE.
20/15=1.33 DRINKS PER MINUTE.
4+2+1.33=7.33 DRINKS PER MINUTE WHEN ALL THREE WORKING.
20/7.33=2.73 MINUTES TO MIS 20 DRINKS WORKING TOGETHER.